OpenStudy (bananas1234):

Find the greatest common factor: 24y^8 + 6y^6

OpenStudy (bananas1234):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the GCF of 24 and 6?

OpenStudy (bananas1234):

6? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

what is the gcf of y^8 and y^6 ?

OpenStudy (bananas1234):

hmm i am not sure... @jim_thompson5910

jimthompson5910 (jim_thompson5910):

well let's think smaller for a sec

jimthompson5910 (jim_thompson5910):

the GCF of y^2 and y^3 is y^2 why? because both y^2 are found in each and this is the largest we can go y^2 = 1*y^2 y^3 = y*y^2 make sense?

OpenStudy (bananas1234):

yes, i think so @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so because y^6 = 1*y^6 y^8 = y^2*y^6 what can we say about the GCF between y^8 and y^6

OpenStudy (bananas1234):

i see

jimthompson5910 (jim_thompson5910):

what is the gcf of y^8 and y^6 ?